# Selection of Filters for Photonic Crystal Spectrometer Using Domain-Aware Evolutionary Algorithms

### v1

@@ 21             Update best-so-far: $\mathbf{x}^{*}\leftarrow\mathbf{y}^{(1)}$;
The global distribution in UMDA-U is likely to have rather a small number of big probabilities. Hence, the candidate solutions sampled from this distribution probably violate the constraint on the number of different filters. Therefore, we extend this algorithm further to overcome this issue.
The idea of Probabilistic Logic Sampling, introduced for Bayesian Networks in [65], is applied in Estimation of Distribution algorithms that address permutations-based COPs [66]. We employ this idea to avoid repetitions of the same filter multiple times. Specifically, we create the conditional distribution $P(\mathrm{x}_{i}|\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1})$ to sample the component $\mathrm{x}_{i}$ of the candidate solution $\mathbf{x}$. If $\mathrm{x}_{i}\in\left\{\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1}\right\}$, then $P(\mathrm{x}_{i}|\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1})$ is set to 00, otherwise it is set to $C\cdot\mathbf{p}_{\mathrm{x}_{i}}$. Constant $C$ is chosen in such a way that: $\sum_{\mathrm{x}_{i}\in\mathbb{L}}P(\mathrm{x}_{i}|\mathrm{x}_{1},\mathrm{x}_{2},\dots,\mathrm{x}_{i-1})=1.$ We call this modification of UMDA-U as UMDA-U-PLS.
When $\lambda$ new points are created, they are ordered in ascending order of the value assigned by $F$ in line 5. Then, the underlining distribution $\mathbf{p}$ is updated. In line 5 we calculate the number of times that filter $j$ occurred in the $\mu$ top-performing candidate solutions. This number is divided by the total number of components in those points. The resulting probability $\mathrm{p}_{j}$ is then adjusted in line 5 to belong to the range $[p_{\text{min}},1-p_{\text{min}}]$, by the following rule:
@@ 5.5 Numerical Results
Reject $H_{0}$? $\mathcal{C}(\bm{x})>12$ $\mathcal{C}(\bm{x})=16$ Reject $H_{0}$? $\mathcal{C}(\bm{x})>12$ $\mathcal{C}(\bm{x})=16$
@@
In this section, we consider the leading algorithms and their distance-driven analogs. That is $(\mu+\lambda)$ EA (Section 4.1.1), $(\mu/2+\lambda)$ EA (Section 4.1.2), UMDA (Section 4.3.2), DD-$(\mu/2+\lambda)$ EA with distances $d_{1},d_{2}$ (Section 5.2), UMDA-U, UMDA-U-PLS-Dist and UMDA-U-PLS-Dist with distances $d_{1},d_{2}$ (Section 5.4). We denote the set of these algorithms as $\mathbb{A}_{2}$. Notably, we did not include DD-$(\mu+\lambda)$ EA to $\mathbb{A}_{2}$ because it was outperformed by DD-$(\mu/2+\lambda)$ EA, which emphasizes the usefulness of the crossover operator on this SCOP.
The results of the application of those algorithms to OFS are presented in Figure 12. From part (b) we see that the two worst algorithms are distance-driven analogs of EA with $d_{1}$ and $d_{2}$. While DD-$(\mu/2+\lambda),d_{1}$ EA manages to find solutions better than $\bm{x}_{\text{base}}$ relatively quickly, DD-$(\mu/2+\lambda),d_{2}$ EA slowly converges to the baseline and then stagnates. Notably, their analog without distance-driven mutation operator outperforms both algorithms at the end of the optimization. However, $(\mu/2+\lambda)$ EA and $(\mu+\lambda)$ EA are slightly worse than DD-$(\mu/2+\lambda),d_{1}$ EA, when very small computational budget is spent $\approx 200$ evaluations of $\widehat{\mathrm{F}}_{K}$. Therefore, a distance-driven analog of EA can produce better solutions when given a tiny computational budget.
@@
One of the possible reasons why distance-driven EAs are worse than conventional EA is big jumps in the search space caused by the relatively huge value of the mean in the distribution over step-sizes $\mathcal{D}.$ It is possible to use parameter-control methods to adjust this parameter during the optimization process, as proposed, for example, in Evolutionary Strategies [67]. Another way is to use a fixed parameter sampled from a heavy-tailed distribution [45].
@@
From part (b) of Figure 12 we see that UMDA-U performs significantly better than UMDA. However, analysis of the best-so-far solutions obtained by UMDA-U shows that each of them has at most 5 different filters, which violates the constraint in Eq. (2b) by far. It is evident from part (a) of Figure 12 that objective values of the solutions produced in different runs of UMDA are significantly different. This causes the rugged behavior of the line that corresponds to UMDA in the considered chart. Therefore, an extension of this algorithm, called UMDA-U-PLS was proposed. UMDA-U-PLS produces solutions with exactly 16 different filters, surpassing UMDA in terms of average best-so-far solution throughout the whole optimization process. This demonstrates that it is faster to learn one distribution over $\mathbb{L}$ than $M$ distributions over $\mathbb{L}$. However, part (a) of Figure 12 depicts jumps in the line that corresponds to UMDA-U-PLS. The jumps happen less frequently than in UMDA-U, however, certain runs still produce relatively bad-performing individuals sampled from the learned conditional distribution.
@@
Notably, UMDA-U-PLS-Dist algorithm with $d_{1}$ performs better than UMDA-U-PLS in terms of the average best-so-far solution, as shown in part (b) of Figure 12. This suggests that the distance-based heuristic used to define the conditional distribution over filters helps to generate better candidate solutions. However, UMDA-U-PLS-Dist algorithm with $d_{2}$ performs worse than its counterpart with $d_{1}$ during the entire optimization process. We can conclude from it that metric $d_{1}$ suits better suited for this type of heuristic as well as for distance-driven EA. At the same time, UMDA-U-PLS-Dist with both distances is robust, as there are no significant jumps in the corresponding lines of Figure 12 (a).
@@
$H_{0}$ is rejected when the pvalue in MWU is smaller than $0.05$. We will say that the vector $\left(\mathrm{g}^{(b)}\!\left[A(\xi^{(i)})\right]\right)_{i=1}^{T}$ is result of the algorithm $A\in\mathbb{A}_{2}$ obtained in $T$ independent runs.
@@ 6 Analysis of the Obtained Solutions to OFS Problem
In this section, we analyze a subset of all candidate solutions $\mathbb{S}$ considered in $n=100$ independent runs by the algorithm UMDA-U-PLS-Dist with $d_{1}$. We now describe how we selected the subset $\mathbb{G}\subset\mathbb{S}$. Remember that the reference solution, described in Section 2, is denoted as $\bm{x}_{\text{base}}$. Considering the points of $\mathbb{S}$ in ascending order of the objective value, that was assigned to them during the optimization process. We select the set $\mathbb{G}$ such that:
@@
Algorithmically we greedily add points from $\mathbb{S}$ to $\mathbb{G}$ in the ascending order of the objective value, that was assigned to the solutions during the optimization process. If at least one of the conditions is not satisfied, then the point is not included in $\mathbb{G}.$ We used constants $D_{\text{min}}=0.05$ and $f_{\text{max}}=\widehat{\mathrm{F}}_{K}\!\left(\bm{x}_{\text{base}}\right)/4\approx 3\cdot 10^{-4},$ where $K=10^{4}.$ According to our experiments, the chosen distance $D_{\text{min}}$ is twice smaller than the distance between a pair of points chosen uniformly at random. Therefore, the set of solutions $\mathbb{G}$ is diverse and high-performing since every element is at least 4 times better than $\bm{x}_{\text{base}}.$ An example of selection $\mathbb{G}$ according to the described procedure is visualized in Figure 14. We will say that the set $\mathbb{G}$ contains final solutions to the OFS problem.
If $|\mathbb{G}|=m$, then let us call the solutions in $\mathbb{G}$ as $\left\{\bm{x}^{(i)}\,\middle|\,i\in\overline{1,m}\right\}$. In our numerical run, we obtained $m=114$ solutions. Following the recommendation in Section 2.8.3 of [22], we reevaluated the objective value for each of them using $K=10^{4}$. We will assume that $\widehat{\mathrm{F}}_{K}\!\left(\bm{x}^{(1)}\right)\leq\widehat{\mathrm{F}}_{K}\!\left(\bm{x}^{(2)}\right)\leq\dots\leq\widehat{\mathrm{F}}_{K}\!\left(\bm{x}^{(m)}\right)$.
In Figure 15 we visualized the objective values of all the found final solutions $\mathbb{G}$. Notably, all of them outperform the baseline solution $\bm{x}_{\text{base}}$ by far. The diversity of the set $\mathbb{G}$ provides practitioners with the flexibility to choose solutions beyond the one with the optimal objective value, denoted by $\bm{x}^{(1)}$. This can be particularly relevant when unforeseen constraints arise after the optimization process. These constraints might include minimizing the manufacturing cost, given that the construction cost is different for filters in $\mathbb{L}$.
@@
In Figure 16 we outline the transmission profiles of every filter in the top 3 final solutions. Surprisingly, it appears that the found solutions contain transmission profiles with little or no sharp features, but more smooth sections with either a low or high transmission. These results show that filters with these features improve the performance of the Trace Gas Measurement Device.
@@
Figure 17: Distribution of relative divination from the ground-truth in one gas retrieval measurement for the baseline solution in (a) and the top-1 found solution in (b). Each random variable is sampled $K=10^{4}$ times and the resulting values are split into 60 segments of equal length. We plot the maximum likelihood Gaussian Distribution that produces the observed samples. In part (b) we copy the Gaussian from part (a) to visualize the stochastic dominance of the solution $\bm{x}^{(1)}$ over $\bm{x}_{\text{base}}.$
In Figure 17 we propose a comparison of the best-found solution against the baseline solution. This analysis shows that the trace gas retrieval algorithm precision follows a Gaussian distribution for both the reference solution (base) as well as the found top solution. The reduced width of the distribution indicates the performance increase of the Trace Gas Measurement Device.
@@ 7 Conclusion
This work investigated the Optimal Filter Selection problem as an instance of the Stochastic Combinatorial Optimization Problem (SCOP). We developed a simulator that mimics the gas retrieval process, including the introduction of artificial noise to the input signal. This noise ultimately translates to the output of the retrieved gas. Our objective was to identify a set of filters that minimizes the expected error after gas retrieval. To achieve this, we proposed a formalization of OFS, which allowed us to define and analyze an objective function that aligns with our intuitive criteria for an optimal solution. Finally, various metaheuristics were employed to generate sets of filters with minimal expected error based on the defined objective function.
@@
We evaluated the performance of various metaheuristics applied to the OFS problem. Based on this analysis, we identified the leading algorithms. To further enhance their efficiency, we developed distance-driven extensions that leverage domain-specific information. This information is encoded as a metric on the search space of all possible OFS solutions. We analyzed and compared the effectiveness of two proposed metrics. Furthermore, to address the internal optimization problem within one of the algorithms, we proposed a novel heuristic for solving the inverse of the Linear Assignment Problem (LAP). We then compared the performance of the distance-driven algorithms against their original counterparts. The comparison was reinforced with statistical testing. This analysis revealed that the proposed algorithm UMDA-U-PLS-Dist with the $d_{1}$ metric stands out as the most efficient and robust solver among the considered approaches.
@@
To analyze the performance of the selected filters, we examined all candidate solutions generated by the algorithm across multiple independent runs. We then identified a diverse subset of high-performing solutions for further analysis and comparison. These findings suggest that filters exhibiting smooth sections (no sharp transitions) with a high contrast in the total transmission lead to an improved performance in the Trace Gas Measurement Device. Furthermore, we observed a significant improvement in the top-performing solution compared to the baseline, indicating the effectiveness of the proposed approach. Additionally, the distribution of trace gas retrieval algorithm precision across different filter sets follows a Gaussian distribution.

### v2

@@ 21      Update best-so-far: $\mathbf{x}^{*}\leftarrow\mathbf{y}^{(1)}$;
When $\lambda$ new points are created, they are ordered in ascending order of the value assigned by $F$ in line 5. Then, the underlining distribution $\mathbf{p}$ is updated. In line 5 we calculate the number of times that filter $j$ occurred in the $\mu$ top-performing candidate solutions. This number is divided by the total number of components in those points. The resulting probability $\mathrm{p}_{j}$ is then adjusted in line 5 to belong to the range $p_{\text{min}},1-p_{\text{min}}]$, by the following rule:
@@ 5.6 Numerical Results
In this section, we consider the leading algorithms and their distance-driven analogs. The algorithms are categorized in Table 2. We group them by their broader classes for clarity. That is $(\mu+\lambda)$ EA (Section 4.1.1), $(\mu/2+\lambda)$ EA (Section 4.1.2), UMDA (Section 4.3.2), DD-$(\mu/2+\lambda)$ EA with distances $d_{1},d_{2}$ (Section 5.3), UMDA-U, UMDA-U-PLS-Dist and UMDA-U-PLS-Dist with distances $d_{1},d_{2}$ (Section 5.5). We denote the set of these algorithms as $\mathbb{A}_{2}$. Notably, we did not include DD-$(\mu+\lambda)$ EA to $\mathbb{A}_{2}$ because it was outperformed by DD-$(\mu/2+\lambda)$ EA, which emphasizes the usefulness of the crossover operator on this SCOP.
Table 2: Algorithms and their legend used in Figure 9
Algorithm Discussed in Legend in Figure 9
Evolutionary Algorithms
$(\mu+\lambda)$ EA Section 4.1.1 [Uncaptioned image]
$(\mu/2+\lambda)$ EA Section 4.1.2 [Uncaptioned image]
UMDA and Extensions
UMDA Section 4.3.2 [Uncaptioned image]
UMDA-U Section 5.5 [Uncaptioned image]
UMDA-U-PLS Section 5.5 [Uncaptioned image]
Distance-Driven Algorithms
$D$-$(\mu/2+\lambda),d_{1}$ EA Section 5.3 [Uncaptioned image]
$D$-$(\mu/2+\lambda),d_{2}$ EA Section 5.3 [Uncaptioned image]
UMDA-U-PLS-Dist, $d_{1}$ Section 5.5 [Uncaptioned image]
UMDA-U-PLS-Dist, $d_{2}$ Section 5.5 [Uncaptioned image]
The results of the application of those algorithms to OFS are presented in Figure 9. Part (a) shows the estimated average quality of the candidate solution proposed at each evaluation step. The degree of short-term variability reflects how the output variations influence the sampling process. When the best-so-far curve looks smooth, but the quality-of-solution curve oscillates, it indicates that the sampling distribution does not stabilize.
@@
From part (b) we see that the two worst algorithms are distance-driven versions of EA with $d_{1}$ and $d_{2}$. While DD-$(\mu/2+\lambda),d_{1}$ EA manages to find solutions better than $\bm{x}_{\text{base}}$ relatively quickly, DD-$(\mu/2+\lambda),d_{2}$ EA slowly converges to the baseline and then stagnates. Notably, their versions without distance-driven mutation operator outperforms both algorithms at the end of the optimization. However, $(\mu/2+\lambda)$ EA and $(\mu+\lambda)$ EA are slightly worse than DD-$(\mu/2+\lambda),d_{1}$ EA, when very small computational budget is spent $\approx 200$ evaluations of $\widehat{\mathrm{F}}_{K}$. Therefore, a distance-driven analog of EA can produce better solutions when given a tiny computational budget.
Reject $H_{0}$? $\mathcal{C}(\bm{x})>12$ ? $\mathcal{C}(\bm{x})=16$ ? Reject $H_{0}$? $\mathcal{C}(\bm{x})>12$ ? $\mathcal{C}(\bm{x})=16$ ?
@@
One of the possible reasons why distance-driven EAs are worse than conventional EA is big jumps in the search space caused by the big value of the mean in the distribution over step-sizes $\mathcal{D}.$ It is possible to use parameter-control methods to adjust this parameter during the optimization process, as proposed, for example, in Evolutionary Strategies [63]. Another way is to use a fixed parameter sampled from a heavy-tailed distribution [41].
@@
From part (b) of Figure 9 we see that UMDA-U performs significantly better than UMDA. However, analysis of the best-so-far solutions obtained by UMDA-U shows that each of them has at most 5 different filters, which violates the constraint in Eq. (1b) by far. It is evident from part (a) of Figure 9 that objective values of the solutions produced in different runs of UMDA are significantly different. This causes the rugged behavior of the line that corresponds to UMDA in the considered chart. Therefore, an extension of this algorithm, called UMDA-U-PLS was proposed. UMDA-U-PLS produces solutions with exactly 16 different filters (this number of filters is used for the reasons discussed in Sec. 1.3), surpassing UMDA in terms of average best-so-far solution throughout the whole optimization process. This demonstrates that it is faster to learn one distribution over $\mathbb{L}$ than $M$ distributions over $\mathbb{L}$. However, part (a) of Figure 9 depicts jumps in the line that corresponds to UMDA-U-PLS. The jumps happen less frequently than in UMDA-U, however, certain runs still produce relatively bad-performing individuals sampled from the learned conditional distribution.
@@
Notably, the improved UMDA-U-PLS-Dist algorithm with $d_{1}$ performs better than UMDA-U-PLS in terms of the average best-so-far solution, as shown in part (b) of Figure 9. This suggests that the distance-based heuristic used to define the conditional distribution over filters helps to generate better candidate solutions. However, UMDA-U-PLS-Dist algorithm with $d_{2}$ performs worse than its counterpart with $d_{1}$ during the entire optimization process. We can conclude from it that metric $d_{1}$ suits better suited for this type of heuristic as well as for distance-driven EA. At the same time, UMDA-U-PLS-Dist with both distances is robust, as there are no significant jumps in the corresponding lines of Figure 9 (a).
@@
$H_{0}$ is rejected when the pvalue in MWU is smaller than $\alpha=0.05$. We will say that the vector $\left(\mathrm{g}^{(b)}\!\left[A(\xi^{(i)})\right]\right)_{i=1}^{T}$ is result of the algorithm $A\in\mathbb{A}_{2}$ obtained in $T$ independent runs.
@@ 6 Analysis of the Obtained Solutions to OFS Problem
In this section, we analyze a subset of all candidate solutions $\mathbb{S}$ considered in $n=100$ independent runs by the algorithm UMDA-U-PLS-Dist with $d_{1}$. We now describe how we selected the subset $\mathbb{G}\subset\mathbb{S}$ to locate a diverse set of high-performing solutions. Remember that the reference solution, described in Section 2, is denoted as $\bm{x}_{\text{base}}$ and we solve the minimization problem. Considering the points of $\mathbb{S}$ in ascending order of the objective value (from worst to best), that was assigned to them during the optimization process. We select the set $\mathbb{G}$ such that:
@@
Algorithmically we greedily add points from $\mathbb{S}$ to $\mathbb{G}$ in the ascending order of the objective value, that was assigned to the solutions during the optimization process. If at least one of the conditions is not satisfied, then the point is not included in $\mathbb{G}.$ We used constants $D_{\text{min}}=0.05$ and $f_{\text{max}}=\widehat{\mathrm{F}}_{K}\!\left(\bm{x}_{\text{base}}\right)/4\approx 3\cdot 10^{-4},$ where $K=10^{4}.$ According to our experiments, the chosen distance $D_{\text{min}}$ is twice smaller than the distance between a pair of points chosen in a series of experiments uniformly at random. Therefore, the set of solutions $\mathbb{G}$ is diverse and high-performing since every element is at least 4 times better than $\bm{x}_{\text{base}}.$ An example of selection $\mathbb{G}$ according to the described procedure is visualized in Figure 11. We will say that the set $\mathbb{G}$ contains final solutions to the OFS problem.
Figure 11, the starts represent solutions found in optimization landscape. Colors depict values of the objective function, with more blue representing better values. Set $\mathbb{G}$ contains a subset of these solutions. Not all blue solutions are selected for this set due to the minimal pairwise distance constraint $D_{\text{min}}$. Also, solutions distant from selected ones not chosen for $\mathbb{G}$ due to their objective value being worse than $f_{\text{max}}$ constraint.
Let us call the solutions in $\mathbb{G}$ as $\left\{\bm{x}^{(i)}\,\middle|\,i\in\overline{1,m}\right\}$. In our numerical run, we obtained $m=|\mathbb{G}|=114$ solutions. Following the recommendation in Section 2.8.3 of [18], we reevaluated the objective value for each of them using $K=10^{4}$. We will assume that $\widehat{\mathrm{F}}_{K}\!\left(\bm{x}^{(1)}\right)\leq\widehat{\mathrm{F}}_{K}\!\left(\bm{x}^{(2)}\right)\leq\dots\leq\widehat{\mathrm{F}}_{K}\!\left(\bm{x}^{(m)}\right)$.
In Figure 12, we visualized the objective values of all the found final solutions $\mathbb{G}$. Notably, all of them outperform the baseline solution $\bm{x}_{\text{base}}$ by far. The diversity of the set $\mathbb{G}$ provides practitioners with the flexibility to choose solutions beyond the one with the optimal objective value, denoted by $\bm{x}^{(1)}$. This can be particularly relevant when unforeseen constraints arise after the optimization process. These constraints might include minimizing the manufacturing cost, given that the construction cost is different for filters in $\mathbb{L}$.
@@
In Figure 13, we outline the transmission profiles of every filter in the top 3 final solutions. Surprisingly, it appears that the found solutions contain transmission profiles with little or no sharp features, but more smooth sections with either a low or high transmission. These results show that filters with these features improve the performance of the Trace Gas Measurement Device.
@@
Figure 14: Distribution of relative deviation from the ground-truth in one gas retrieval measurement for the baseline solution in (a) and the top-1 found solution in (b). Each random variable is sampled $K=10^{4}$ times and the resulting values are split into 60 segments of equal length. We plot the maximum likelihood Gaussian Distribution that produces the observed samples. In part (b) we copy the Gaussian from part (a) to visualize the stochastic dominance of the solution $\bm{x}^{(1)}$ over $\bm{x}_{\text{base}}.$
In Figure 14, we propose a comparison of the best-found solution against the baseline solution. This analysis shows that the trace gas retrieval algorithm precision follows a Gaussian distribution for both the reference solution (base) as well as the found top solution. The reduced width of the distribution indicates an increase in the performance of the Trace Gas Measurement Device.
@@ 7 Conclusion
This work investigated the Optimal Filter Selection (OFS) problem as an instance of the Stochastic Combinatorial Optimization Problem (SCOP). We developed a simulator that mimics the gas retrieval process, including the introduction of artificial noise to the input signal, which translates to the output of the retrieved gas. Our objective was to identify a set of filters that minimizes the expected error after gas retrieval. To achieve this, we proposed a formalization of OFS, which allowed us to define and analyze an objective function that aligns with our intuitive criteria for an optimal solution. Finally, various metaheuristics were employed to generate sets of filters with minimal expected error based on the defined objective function.
@@
We evaluated the performance of various metaheuristics applied to the OFS problem. Based on this analysis, we identified the leading algorithms. To further enhance their efficiency, we developed distance-driven extensions that leverage domain-specific information. This information is encoded as a metric on the search space of all possible OFS solutions. We analyzed and compared the effectiveness of two proposed metrics. Furthermore, to address the internal optimization problem within one of the algorithms, we proposed a novel heuristic, called Distance‑Driven Assignment Evolutionary Algorithm (DDA‑EA), for solving the inverse of the Linear Assignment Problem (LAP). We then compared the performance of the distance-driven algorithms against their original counterparts. The comparison was reinforced with statistical testing. This analysis revealed that the proposed first-of-its-kind algorithm UMDA-U-PLS-Dist with the $d_{1}$ metric stands out as the most efficient and robust solver among the considered approaches. The code and tools used to simulate trace gas measurement devices, apply optimization algorithms for filter selection, and perform the experiments described in this work can be accessed at https://github.com/AntKirill/CompressSpecLIACS-Public.
@@
To analyze the performance of the selected filters, we examined all candidate solutions generated by the algorithm across multiple independent runs. We then identified a diverse subset of high-performing solutions for further analysis and comparison. These findings suggest that filters exhibiting smooth sections (no sharp transitions) with a large local difference in the transmission lead to improved performance in the Trace Gas Measurement Device. Furthermore, we observed a significant improvement in the top-performing solution compared to the baseline, indicating the effectiveness of the proposed approach. Additionally, the distribution of trace gas retrieval algorithm precision across different filter sets follows a Gaussian distribution.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors utilized ChatGPT to correct grammatical errors and rephrase some sentences to enhance clarity. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Appendix A Properties of the OFS Objective Function
Let us show the properties of the OFS objective function defined in Eq. (3). First of all, it conforms to our intuitive criteria 1 for the assessment of filter selections. Indeed, for the fixed solutions $\bm{x},\bm{y}\in\mathscr{L}$ the requirement can be rewritten as follows:
$\operatorname{E}^{2}\left[\mathrm{D}(\bm{x},\xi)\right]\leq\operatorname{E}^{2}\left[\mathrm{D}(\bm{y},\xi)\right]\operatorname{Var}\left[\mathrm{D}(\bm{x},\xi)\right]<\operatorname{Var}\left[\mathrm{D}(\bm{y},\xi)\right].$ (12)
Let us demonstrate how function $F$ influences the additional difference in means needed to compensate for the increase in variances. Consider some fixed selections of filters $\bm{x},\bm{y}\in\mathscr{L}$ that satisfy Eq. (13). Due to the symmetry between $\bm{x}$ and $\bm{y}$, we assume without loss of generality that:
$\mathrm{F}(\bm{x})\leq\mathrm{F}(\bm{y})$ (14)
Case 1)
Case 2)
$\dfrac{\operatorname{E}^{2}(\mathrm{Y})}{\operatorname{Var}(\mathrm{Y})}\cdot\left(\operatorname{E}^{2}(\mathrm{Y})-\operatorname{E}^{2}(\mathrm{X})\right)\geq\dfrac{\operatorname{E}^{2}(\mathrm{Y})}{\operatorname{Var}(\mathrm{Y})}\cdot\left(\operatorname{Var}(\mathrm{X})-\operatorname{Var}(\mathrm{Y})\right).$ (16)
Refer to caption Refer to caption
Case 3)
Analysis.
We compute the sample mean of the set $\mathbb{V}$, which gives an approximation of the value in LHS of Eq. (17) and Eq. (18) expected while optimizing the function $F$. The $\ln$ of this value equals 6.6-6.6, and it is depicted by the gray plane in Figure 15. The part of the surface that is lower than the gray plane in Figure 15 depicts the set of ratios of expected values and means that lead to grading $\mathrm{X}$ better than $\mathrm{Y}$ in average over the set $\mathbb{V}$. Conversely, the region above the plane represents ratios leading to worse average grading of $\mathrm{X}$ compared to $\mathrm{Y}$.
When $\operatorname{Var}(\mathrm{X})$ grows relative to $\operatorname{Var}(\mathrm{Y})$, the value ${\operatorname{Var}(\mathrm{X})}/{\operatorname{Var}(\mathrm{Y})}$ becomes bigger. As demonstrated in Figure 15, in this situation it becomes less likely to assess $\mathrm{X}$ better than $\mathrm{Y}$. The intersection between the surface and the gray plane in Figure 15 occurs when the ratio of variances, $\operatorname{Var}(\mathrm{X})/\operatorname{Var}(\mathrm{Y})$, is less than $1.0024$. This intersection depicts the maximum tolerable ratio in variances where $\mathrm{X}$ can still be graded better than $\mathrm{Y}$ by the function $F$, on average over the set $\mathbb{V}$. When such grading is possible, we see that it becomes less likely when values of $\operatorname{E}(\mathrm{X})$ and $\operatorname{E}(\mathrm{Y})$ are closer in absolute value. Moreover, it is clear that for the fixed ratio of expected value the RHS of Eq. (17) is monotonically increasing, so grading of $\mathrm{X}$ better than $\mathrm{Y}$ under the assumption in Eq. (13) is possible only when variances are not too different but means compensate this difference by far. In summary, we demonstrated how the means must adjust to offset the rise in variances. Although the results of this analysis are not directly applied in the subsequent sections of the paper, they remain crucial as they provide guarantees regarding the candidate solutions we favor.
Appendix B Metric space for OFS
We remind the following classical definition of the metric space in Definition 3.
Definition 3 (Metric Space).
  1. 1.
    $d(x,y)\geq 0$ and $d(x,y)=0\iff x=y$,
  2. 2.
    $d(x,y)=d(y,x)$,
  3. 3.
    $d(x,z)\leq d(x,y)+d(y,z)$.
The function $d$ is called the metric on $X$.
Proposition 1.
Proof.
  1. 1.
  2. 2.
  3. 3.
    There exist bijections $\pi_{1},\pi_{2}\in\mathbb{P}$ such that:
    $\mathscr{D}(\bm{x},\bm{y})+\mathscr{D}(\bm{y},\bm{z})$
    $\geq\sum_{i=1}^{k}d_{\mathbb{L}}(x_{i},z_{\pi_{2}\circ\pi_{1}(i)})$